OpenStudy (anonymous):

A student scores 72 out of 100 on a test. If the maximum score on the next test is also 100 points, what score does the student need to maintain at least an average of 80?

OpenStudy (anonymous):

let the score be x then (72+x)/2>=80

OpenStudy (anonymous):

does it help?

OpenStudy (anonymous):

The 2 tests' points add together to make 200. A percentage is in the form \[\ \frac{x}{100} \], So divide our 200 score by 2 to get it in 100-terms Call y the score on the 2nd test. We want \[\frac{72+y}{2}\] to equal at least 80 \[\frac{72+y}{2} >/= 79\]

OpenStudy (anonymous):

can you help me through it

OpenStudy (anonymous):

Re-arrange and find y

OpenStudy (anonymous):

@henpen why 79

OpenStudy (anonymous):

not 80

OpenStudy (anonymous):

Sorry, 80, the 'or equal to' sign was breaking on me!